Julie travels 6 times faster than Eric. Traveling In opposite directions, they are 196 miles apart after 2 hours. Find their rate of travel
Answer by Alan3354(69443)

Julie travels 6 times faster than Eric. Traveling In opposite directions, they are 196 miles apart after 2 hours. Find their rate of travel
----------------
6 times faster = 7 times as fast
---
E = Eric's speed
J = Julie's speed
----
d = r*t
r = d/t = 196/2
r = 98 mi/hr
-----
E + J = 98
J = 7E
---
E + 7E = 98
8E = 98
E = 12.25 mi/hr
J = 85.75 mi/hr